# An Early Look at the Performance of IGRINS–2 at Gemini-North with Application to the ultrahot Jupiter, WASP-33 b

### v1

@@ Abstract
Ground-based high-resolution spectroscopy enables precise molecular detections, and velocity-resolved atmospheric dynamics, offering a distinct advantage over low-resolution methods for exoplanetary atmospheric studies. IGRINS-2, the successor to IGRINS, features improved throughput and enhanced sensitivity to carbon monoxide by shifting its K-band coverage by 36 nm to longer wavelengths. To evaluate its performance, we attempt to investigate the atmospheric characteristics of WASP-33 b. Observations were conducted on 2024 January 7 for a total of 2.43 hours; This includes 1.46 hours in the pre-eclipse phase to capture the planet’s thermal emission spectrum. Our analysis, employing a composite atmospheric model, confirms the presence of a thermally inverted atmosphere on the planet’s dayside with a signal-to-noise ratio (SNR) of 7.5. More specifically, we capture CO, H2O, and OH with SNRs of 5.2, 4.2, and 4.4, respectively. These results are consistent with previous studies and demonstrate that IGRINS-2 is well-suited for detailed investigation of exoplanetary atmospheres. We anticipate that future observations with IGRINS-2 will further advance our understanding of exoplanetary atmospheres.
@@ 1 Introduction
High-resolution infrared spectroscopy (R 15,000) has become a cornerstone for characterizing exoplanetary atmospheres, offering unique advantages over low-resolution spectroscopy. Unlike low-resolution spectroscopy, which primarily probes deeper layers of an atmosphere, high-resolution spectroscopy enables the detection of molecular species across a broader range of pressures, providing insights into thermal profiles and atmospheric dynamics (Snellen et al., 2010). Additionally, high-resolution spectroscopy is highly sensitive to velocity-resolved information, allowing for the study of wind patterns and rotational dynamics (Wardenier et al., 2024). These capabilities offer a deeper understanding of atmospheric chemistry and planetary formation histories (Öberg et al., 2011; Booth et al., 2017; Chachan et al., 2023). High-resolution cross-correlation spectroscopy (HRCCS), in particular, has become an essential tool for exploring exoplanetary atmospheres. By aligning observed spectra with theoretical templates, HRCCS allows for the detection of molecular and atomic species and provides insights into thermal and velocity structures. This technique has been successfully applied to various orbital geometries (i.e. transmission and thermal emission), enabling a comprehensive investigation of exoplanetary atmospheres (Snellen et al., 2010; de Kok et al., 2013; Brogi et al., 2023).
@@
Previous high-resolution spectroscopy studies have revealed a thermally inverted atmosphere on the dayside of WASP-33 b, characterized by significant photodissociation of H2O and the detection of various molecules and atomic species. Emission spectroscopy has identified CO, OH, and Fe I as prominent components in the dayside atmosphere, while transmission spectroscopy has detected AlO and Si I in the terminator region (Nugroho et al., 2021; Cont et al., 2021, 2022a, 2022b; Yan et al., 2022; van Sluijs et al., 2023a). These findings suggest a complex atmospheric structure, with early inferences pointing to a thermal inversion on the dayside. Yang et al. (2024) extend this understanding by suggesting that thermal inversion also occurs in the terminator region, offering new insights into the atmospheric dynamics and chemical processes of WASP-33 b.
The recent study by Finnerty et al. (2023, hereafter, F23) using the Keck Planet Imager and Characterizer (KPIC, R 35,000) provided comprehensive insights into the atmospheric composition of WASP-33 b. Their analysis validated a thermally inverted dayside atmosphere with significant detections of CO, H2O, and OH in emissions. Following the detection of molecular species, Bayesian atmospheric retrievals (Brogi & Line, 2019) were performed to place quantitative constraints on the composition of WASP-33 b’s dayside atmosphere. The model fits indicated a carbon- and metal-enriched environment, with a gas-phase C/O ratio of 0.80.2 and evidence for a metal-enriched atmosphere (2–15 solar). These findings are consistent with formation scenarios that involve accretion near the CO or CO2 snow lines (Öberg et al., 2011) or enrichment by carbon-rich grains interior to the H2O snowline (Chachan et al., 2023), as suggested by F23. Additionally, F23 found tentative evidence for a 12CO/13CO ratio of 50 consistent with values expected in protoplanetary disks.
@@ 2 Upgrades on IGRINS-2
The upgrades from IGRINS to IGRINS-2 include enhanced sensitivity and a shift in spectral coverage, which are critical for HRCCS of exoplanetary atmospheres. This shift resulted from the process of aligning the echellogram on the detector by adjusting the main disperser, the immersion grating. The shift toward longer wavelengths was not intentional; rather, the wavelength range was determined while optimizing optical performance, avoiding bad pixels, etc., within the required K-band wavelength range of 1.96–-2.46 µm. To isolate instrumental differences from site-specific effects (e.g., telescope throughput, atmospheric transmission), we compared flat lamp frames from the calibration systems of both instruments. The major changes in optics include replacing the mirror of the IGRINS’s calibration system with a beam splitter in IGRINS-2 and using a dichroic mirror with enhanced response at the edges of each channel in the spectrograph. We corrected these differences by interpolating and normalizing the response function based on the transmittance of each optical component, though these corrections had an insignificant impact on the final spectral response (see Figure 1).
@@ 2.1 K-Band Spectral Shift
One of the most significant changes in IGRINS-2 is the shifts of the H and K-band spectra to longer wavelengths. Specifically, the K band now extends approximately 36 nm further to longer wavelengths compared to IGRINS. This shift could be crucial because it captures additional spectral features, particularly CO molecular lines that begin to appear around 2.3 µm. By covering more CO lines in the K band, as shown in the bottom right panel of Figure 1, IGRINS-2 improves the detection and characterization of molecular features in exoplanetary atmospheres, resulting in an increased signal in the cross-correlation analysis.
@@ 2.2 Sensitivity Enhancements
Another significant advancement in IGRINS-2 is the improvement in the quantum efficiency (QE) of the detector in the K band. The QE of IGRINS in the K band was approximately 80%–85% (Yuk et al., 2010), whereas IGRINS-2 achieved an average QE of 96% (C. Park et al. in preparation). This increase in QE results in more efficient photon collection, particularly at longer wavelengths, enhancing the overall sensitivity of the instrument and improving the detection of weaker molecular features.
@@ 3.1 Observations
Figure 2 shows the SNR, seeing, and airmass variations during the observations. Throughout the observation period, the relative humidity remained stable at approximately 9%, as recorded by the Maunakea Weather Center111http://mkwc.ifa.hawaii.edu/archive/ at the Canada–France–Hawaii Telescope station, which is located near Gemini-North. Seeing, obtained from the Differential Image Motion Monitor (DIMM), varied between 0″.52 and 0″.97, with an average seeing of approximately 0″.68.
@@ 3.2 Data reduction
The data reduction was performed using the IGRINS Pipeline Package Version 3.0 (PLP; Kaplan et al., 2024), which was initially created for the original IGRINS. The igrins2-dev branch of the pipeline, which contains the necessary modifications for IGRINS-2, was utilized. Flat-fielding was applied to the raw data to adjust for pixel-to-pixel differences. Bad pixels were identified and removed from both flat-off and flat-on images. One-dimensional spectra were extracted using a modified version of Horne (1986) optimal extraction techniques. The final result consists of one-dimensional spectra for each diffraction order.
@@
After initial data reduction, we used the IGRINStransit (Mansfield & Lee, 2024) package for secondary wavelength calibration and barycentric-velocity correction which was described in Line et al. (2021). IGRINStransit also includes routines for the data cube production, pixel trimming, and exclusion of bad orders. We trimmed 100 pixels from the edges of each order due to their low SNRs and discarded bad orders at the edge of each band that were strongly contaminated by telluric absorption lines during calibration. As a result, five orders were discarded out of the total 53 orders: the longest order (2.520–2.554 µm) and the shortest three orders (1.866–1.932 µm) in the K band, as well as the shortest order (1.441–1.461 µm) in the H band. Each remaining order contains 1848 pixels.
To minimize noise from stellar variability and systematics, we cropped the data following the secondary eclipse phase before analysis. The complete dataset (including eclipse observations) was preserved for trail plot visualization to maintain full orbital phase context.
@@
The data cube, produced by stacking the time series spectra for each order and pixel, contains a mix of the stellar spectrum, telluric absorption, planet signal, and photon noise. To extract the planet signal, it is crucial to isolate and remove the stationary stellar and telluric lines, which remain quasi-stationary during observations, while the planet’s spectrum shifts due to its orbital motion. This can be achieved through singular value decomposition (SVD), which decomposes the time-series data into principal components (see Birkby et al., 2013; de Kok et al., 2013; Giacobbe et al., 2021; Line et al., 2021). The dominant stationary features are removed, leaving the planet signal in the residuals. The optimal number of SVD components to remove can vary depending on observing conditions and the level of telluric absorption. Following the approach of F23, we found that using four singular values (SVs) yielded the highest SNR via a cross-correlation, though using three or five SVs resulted in minimal changes within one .
3.3 Wavelength calibration stability
Wavelength calibration stability is critical to the success of HRCCS, as even minor instabilities in the wavelength solution can adversely affect the alignment between observed spectra and atmospheric models during the cross-correlation process. We evaluated the wavelength calibration stability using two steps: First, due to the broad stellar absorption lines caused by the rapid rotation of the host star, the telluric lines were primarily used to evaluate the stability of the initial wavelength solution. Then, a secondary wavelength calibration using scipy.curve_fit was applied to refine the wavelength solution. This corrected any residual shifts in the wavelength calibration across the spectral orders. Figure 3 shows the resulting wavelength calibration shifts over time for both the H & K bands, where each point represents the shift for an individual spectral order.
The analysis demonstrates that these two calibration steps achieved high stability, with wavelength shifts maintained within the sub-pixel range throughout the observations. This level of stability ensures that the Doppler shifts due to planetary motion can be accurately captured during the HRCCS process, minimizing errors in molecular signal detection.
@@ 4.1 Atmospheric models
To perform cross-correlation analysis, we generated thermal emission spectra for WASP-33 b with a resolving power of 250,000 using the petitRADTRANS package (Mollière et al., 2019, 2020). The atmospheric structure is modeled across a pressure range from 10-6 to 102 bar, divided into 100 layers with uniform spacing in logarithmic pressure. Models for each molecule and pressure-temperature (P-T) profile used in the analysis are shown in Figure 4. Consistent with several previous studies, we assumed a thermal inversion profile for the dayside atmosphere (Nugroho et al., 2021; van Sluijs et al., 2023b; Yan et al., 2022; Finnerty et al., 2023) and we adopted Guillot P-T profile (Guillot, 2010). Molecular species included in the models were H2O (Rothman et al., 2010), CO (Li et al., 2015), and OH (Gordon et al., 2022) from HITRAN/HITEMP, chosen for following the results from F23. These models were built based on key parameters from F23, including reference pressure, mass mixing ratios (MMR), and the Guillot P-T profile parameters , and IR (Finnerty et al., 2023, private communication). Stellar and planetary properties were adopted from Chakrabarty & Sengupta (2019), and a PHOENIX model was used for the stellar template (Husser et al., 2013). The parameters used for the atmospheric models are summarized in Table 1.
@@ 4.2 Cross-correlation maps
To validate the performance of IGRINS-2 on HRCCS, we cross-correlated the post-SVD residual spectra with atmospheric models. These models were convolved with the IGRINS-2 instrumental profile to match the resolving power of the observations (). The cross-correlation function (CCF) was computed by systematically shifting the model spectra across a range of radial velocities and calculating the correlation at each step. We followed the same approach described in Brogi & Line (2019) and Line et al. (2021) that is used to detect the atmospheric signal using IGRINS (see Brogi et al., 2023; Kanumalla et al., 2024; Smith et al., 2024; Weiner Mansfield et al., 2024). Frames taken during the secondary eclipse, when the planet is hidden behind its host star, were not included when generating - maps.
@@
Figure 5 presents the resulting cross-correlation maps, generated using the composite atmospheric model with CO, H2O, and OH. A planetary trail before ingress is clearly shown in the top panel. The dashed red line corresponds to the Doppler shift induced by the planet’s orbital motion. The alignment of this trail with the predicted velocity curve, derived from WASP-33 b’s orbital parameters, confirms the detection of the planet’s atmospheric signal. The cross-correlation map (bottom panel in Fig. 5) shows the composite atmospheric thermal emission with SNR = 7.5. The dashed white lines indicate the expected planet’s radial velocity () from F23 and a system velocity () from Gontcharov (2006). SNR is calculated following the same approach outlined in Weiner Mansfield et al. (2024) using the 3 clipping method. The positive cross-correlation value further suggests that the detected spectral lines are in emission rather than absorption, supporting the presence of a thermally inverted dayside atmosphere.
@@
We also attempted to detect individual molecular signals from CO, H2O, and OH in the atmosphere using models in Figure 4. These results are shown in Figure 6, formatted similarly to the result of the composite atmospheric model. CO is robustly detected at the expected planetary velocity, consistent with previous studies and confirming its presence in the planet’s dayside emission. The cross-correlation results of H2O and OH exhibit velocity shifts, which may suggest additional atmospheric processes at play, as individual molecular species have shown similar shifts in previous studies (Brogi et al., 2023; Cont et al., 2021, 2024). One possibility is that the intense dayside heating leads to the thermal dissociation of H2O into H and OH, resulting in non-uniform molecular distributions across the atmosphere (e.g., Lothringer et al., 2018; Parmentier et al., 2018; Nugroho et al., 2021, Finnerty et al., in review). Alternatively, strong atmospheric dynamics, such as high-altitude winds or circulation patterns, may contribute to the observed Doppler shifts by redistributing molecular species across different regions of the atmosphere. Follow-up observations, particularly those covering a broader orbital phase range, will be essential to confirm these velocity shifts and further investigate the planet’s atmospheric circulation and chemistry.
Notably, we achieved high SNR molecular detections on the planetary signal (0.05 orbital phase) with an observation time of only 1.46 hours. This is shorter than the typical observation time of 2–4 hours required in similar studies conducted with 8-m class telescopes (Nugroho et al., 2021; Line et al., 2021; Brogi et al., 2023; Finnerty et al., 2023). This result highlights the sensitivity and efficiency of IGRINS-2 for HRCCS.

### v3

# An Early Look at the Performance of IGRINS-2 at Gemini-North with Application to the ultrahot Jupiter, WASP-33 b

@@ Abstract
Ground-based high-resolution spectroscopy enables precise molecular detections and velocity-resolved atmospheric dynamics, offering a distinct advantage over low-resolution methods for exoplanetary atmospheric studies. IGRINS-2, the successor to IGRINS, features improved throughput and enhanced sensitivity to carbon monoxide by shifting its K-band coverage by 36 nm to longer wavelengths. IGRINS is a near-infrared high-resolution spectrograph mounted at McDonald, Lowell, and Gemini-South observatories. Our order-drop test shows this added range improves the CO cross-correlation signal-to-noise ratio (SNR) by 2–-3, confirming a measurable but modest sensitivity gain. To evaluate its performance, we attempt to investigate the atmospheric characteristics of WASP-33 b. Observations were conducted on 2024 January 7 for a total of 2.43 hours; This includes 1.46 hours in the pre-eclipse phase to capture the planet’s thermal emission spectrum. We successfully detect clear cross-correlation signals from molecular species in the dayside atmosphere of WASP-33 b with a combined SNR of 7.4. More specifically, we capture CO, H2O, and OH with SNRs of 6.3, 4.7, and 4.2, respectively. These results are consistent with previous studies and demonstrate that IGRINS-2 is well-suited for detailed investigation of exoplanetary atmospheres. We anticipate that future observations with IGRINS-2 will further advance our understanding of exoplanetary atmospheres.
@@ 1 Introduction
High-resolution infrared spectroscopy (R 15,000) has become a cornerstone for characterizing exoplanetary atmospheres, offering unique advantages over low-resolution spectroscopy. Unlike low-resolution spectroscopy, which primarily probes higher layers of an atmosphere, high-resolution spectroscopy enables the detection of molecular species across a broader range of pressures, providing insights into thermal profiles and atmospheric dynamics (Snellen et al., 2010). Additionally, high-resolution spectroscopy is highly sensitive to velocity-resolved information, allowing for the study of wind patterns and rotational dynamics (Wardenier et al., 2024). These capabilities offer a deeper understanding of atmospheric chemistry and planetary formation histories (Öberg et al., 2011; Booth et al., 2017; Chachan et al., 2023). High-resolution cross-correlation spectroscopy (HRCCS), in particular, has become an essential tool for exploring exoplanetary atmospheres. By aligning observed spectra with theoretical templates, HRCCS allows for the detection of molecular and atomic species and provides insights into thermal and velocity structures. This technique has been successfully applied to various orbital geometries (i.e., transmission and thermal emission), enabling a comprehensive investigation of exoplanetary atmospheres (Snellen et al., 2010; de Kok et al., 2013; Brogi et al., 2023).
@@
Previous high-resolution spectroscopy studies have revealed a thermally inverted atmosphere on the dayside of WASP-33 b, characterized by significant photodissociation of H2O and the detection of various molecules and atomic species. Emission spectroscopy has identified CO, OH, TiO, Fe I, Si I, Ti I, V I as prominent components in the dayside atmosphere (Nugroho et al., 2017, 2020, 2021; Cont et al., 2021, 2022a, 2022b; Yan et al., 2022; van Sluijs et al., 2023a), while transmission spectroscopy has detected hydrogen Balmer lines and Ca II in the terminator region (Yan et al., 2019, 2021). These findings suggest a complex atmospheric structure, with early inferences pointing to a thermal inversion on the dayside. Yang et al. (2024) extend this understanding by suggesting that thermal inversion also occurs in the terminator region, offering new insights into the atmospheric dynamics and chemical processes of WASP-33 b.
The recent study by Finnerty et al. (2023, hereafter, F23) using the Keck Planet Imager and Characterizer (KPIC; R 35,000) provided comprehensive insights into the atmospheric composition of WASP-33 b. Their analysis validated a thermally inverted dayside atmosphere with significant detections of CO, H2O, and OH in emissions. Following the detection of molecular species, Bayesian atmospheric retrievals (Brogi & Line, 2019) were performed to place quantitative constraints on the composition of WASP-33 b’s dayside atmosphere. The model fits indicated a carbon- and metal-enriched environment, with a gas-phase C/O ratio of 0.80.2 and evidence for a metal-enriched atmosphere (2–15 solar). These findings are consistent with formation scenarios that involve accretion near the CO or CO2 snow lines (Öberg et al., 2011) or enrichment by carbon-rich grains interior to the H2O snowline (Chachan et al., 2023), as suggested by F23. Additionally, F23 found tentative evidence for a 12CO/13CO ratio of 50 consistent with values expected in protoplanetary disks.
@@ 2 Upgrades on IGRINS-2
The upgrades from IGRINS to IGRINS-2 include enhanced sensitivity and a shift in spectral coverage, which are critical for HRCCS of exoplanetary atmospheres. This shift resulted from the process of aligning the echellogram on the detector by adjusting the main disperser, the immersion grating. The shift toward longer wavelengths was not deliberate but arose naturally during the optical alignment and optimization process of the spectrograph (Oh et al., 2024). Specifically, physical adjustments were made to the spectrograph’s main disperser, the immersion grating, and the detector to optimize optical performance within the required K-band range of 1.96–2.46 µm. As part of this process, additional adjustments were implemented to minimize the impact of detector artifacts, such as bad pixels, on the spectra. This technical optimization resulted in a slightly shifted coverage at longer wavelengths in K-band. To isolate instrumental differences from site-specific effects (e.g., telescope throughput, atmospheric transmission), we compared flat lamp frames from the calibration systems of both instruments. The major changes in optics include replacing the mirror of the IGRINS’s calibration system with a beam splitter in IGRINS-2 and using a dichroic mirror with enhanced response at the edges of each channel in the spectrograph. We corrected these differences by interpolating and normalizing the response function based on the transmittance of each optical component, though these corrections had an insignificant impact on the final spectral response (see Figure 1).
@@ 2.1 K-Band Spectral Shift
One of the most significant changes in IGRINS-2 is the shifts of the H and K-band spectra to longer wavelengths. Specifically, the K band now extends approximately 36 nm further to longer wavelengths compared to IGRINS. This shift could be crucial because it captures additional spectral features, particularly CO molecular lines that begin to appear around 2.3 µm. By covering more CO lines in the K band, as shown in the bottom right panel of Figure 1, IGRINS-2 improves the detection and characterization of molecular features in exoplanetary atmospheres, resulting in an increased signal in the cross-correlation analysis. The impact of including this extended spectral range on CO, particularly the 71st order (2.48–2.52 µm), which is near the longest spectral order, is tested and discussed further in Section 4.2.
@@ 2.2 Sensitivity Enhancements
Another significant advancement in IGRINS-2 is the improvement in the quantum efficiency (QE) of the detector in the K band. The QE of IGRINS in the K band was approximately 80–85% (Yuk et al., 2010), whereas IGRINS-2 achieved an average QE of 96% (C. Park et al. in preparation). This increase in QE results in more efficient photon collection, particularly at longer wavelengths, enhancing the overall sensitivity of the instrument and improving the detection of weaker molecular features.
@@ 3.1 Observations
Figure 2 shows the SNR, seeing, and airmass variations during the observations. Throughout the observation period, the relative humidity remained stable at approximately 9%, as recorded by the Maunakea Weather Center222http://mkwc.ifa.hawaii.edu/archive/ at the Canada-France-Hawaii Telescope station, which is located near Gemini-North. Seeing, obtained from the Differential Image Motion Monitor (DIMM), varied between and , with an average seeing of approximately .
@@ 3.2 Data reduction
The data reduction was performed using the IGRINS Pipeline Package Version 3.0 (PLP; Kaplan et al., 2024), which was initially created for the original IGRINS. The igrins2-dev branch333https://github.com/igrins/plp/tree/igrins2-dev of the pipeline, which contains the necessary modifications for IGRINS-2, was utilized. Flat-fielding was applied to the raw data to adjust for pixel-to-pixel differences. Bad pixels were identified and removed from both flat-off and flat-on images. One-dimensional spectra were extracted using a modified version of Horne (1986) optimal extraction techniques. The final result consists of one-dimensional spectra for each diffraction order.
@@
After initial data reduction, we used the IGRINS_transit444https://github.com/meganmansfield/IGRINS_transit (Mansfield & Lee, 2024) package for secondary wavelength calibration and barycentric-velocity correction which was described in Line et al. (2021). IGRINStransit also includes routines for the data cube production, pixel trimming, and exclusion of bad orders. We trimmed 100 pixels from the edges of each order due to their low SNRs and discarded orders at the edges of each band that had low transmittance and strong telluric contamination during calibration. As a result, five orders were discarded out of the total 53 orders: the longest order (2.520–2.554 µm) and the shortest three orders (1.866–1.932 µm) in the K band, as well as the shortest order (1.441–1.461 µm) in the H band. Each remaining order contains 1848 pixels. Then, a secondary wavelength calibration using scipy.curve_fit was applied to refine the wavelength solution. This corrected any residual shifts in the wavelength calibration across the spectral orders. Figure 3 shows the resulting wavelength calibration shifts over time for both the H & K bands, where each point represents the shift for an individual spectral order.
To minimize noise contributions from stellar variability and instrumental systematics, we excluded data obtained after the start of the secondary eclipse from subsequent analyses. Specifically, spectra taken during and after the eclipse ingress were removed. While PCA-based methods like SVD typically benefit from larger datasets, including spectra with no expected planetary signal (e.g., those during eclipse) can introduce non-planetary variance into the analysis, thereby suppressing the planetary signal in the decomposition. Excluding these data thus effectively reduces irrelevant variability, enhancing detection significance. The complete dataset (including eclipse observations) was preserved for trail plot visualization to maintain full orbital phase context.
@@
The data cube, produced by stacking the time series spectra for each order and pixel, contains a mix of the stellar spectrum, telluric absorption, planet signal, and photon noise. To extract the planet signal, it is crucial to isolate and remove the stationary stellar and telluric lines, which remain quasi-stationary during observations, while the planet’s spectrum shifts due to its orbital motion. This can be achieved through singular value decomposition (SVD), which decomposes the time-series data into principal components (see Birkby et al., 2013; de Kok et al., 2013; Giacobbe et al., 2021; Line et al., 2021). The dominant stationary features are removed, leaving the planet signal in the residuals. The optimal number of SVD components to remove can vary depending on observing conditions and the level of telluric absorption. Following the approach described by F23, we performed cross-correlation analyses using different numbers of singular values (SVs). We found that using four SVs resulted in the highest SNR. However, the derived signals were robust to small changes in SV selection; specifically, the resulting SNR varied minimally (within one sigma) when using three to five SVs, confirming the stability of our detection.
@@ 4.1 Atmospheric models
To perform cross-correlation analysis, we generated thermal emission spectra for WASP-33 b with a resolving power of 250,000 using the petitRADTRANS package (Mollière et al., 2019, 2020). The atmospheric structure is modeled across a pressure range from 10-6 to 102 bar, divided into 100 layers with uniform spacing in logarithmic pressure. Models for each molecule and pressure-temperature (P-T) profile used in the analysis are shown in Figure 4. Consistent with several previous studies, we assumed a thermal inversion profile for the dayside atmosphere (Nugroho et al., 2021; van Sluijs et al., 2023b; Yan et al., 2022; Finnerty et al., 2023) and we adopted Guillot P-T profile (Guillot, 2010). Molecular species included in the models were H2O (Rothman et al., 2010), CO (Li et al., 2015), and OH (Gordon et al., 2022) from HITRAN/HITEMP, chosen for following the results from F23. These models were built based on key parameters from F23, including reference pressure, mass mixing ratios (MMRs), and the Guillot P-T profile parameters , and IR (Finnerty et al., 2023, private communication). Stellar and planetary properties were adopted from Chakrabarty & Sengupta (2019), and a PHOENIX model was used for the stellar template (Husser et al., 2013). The stellar template is then smoothed to approximate the stellar continuum. The parameters used for the atmospheric models are summarized in Table 1.
@@ 4.2 Cross-correlation maps
To validate the performance of IGRINS-2 with HRCCS, we cross-correlated the post-SVD residual spectra with atmospheric models using the Pearson correlation coefficient. These normalized correlation coefficients measure how well the variations in the observed spectra match those of the shifted model spectra. The atmospheric models were convolved with the IGRINS-2 instrumental profile to match the resolving power of the observations (). The cross-correlation function (CCF) was computed by systematically shifting the model spectra across a range of radial velocities and calculating the correlation at each step. We followed the same approach described in Brogi & Line (2019) and Line et al. (2021) that is used to detect the atmospheric signal using IGRINS (see Brogi et al., 2023; Kanumalla et al., 2024; Smith et al., 2024; Weiner Mansfield et al., 2024). Frames taken during the secondary eclipse, when the planet is hidden behind its host star, were not included when generating - maps.
@@
Figure 5 presents the resulting cross-correlation maps, generated using the composite atmospheric model with CO, H2O, and OH. A planetary trail before ingress is clearly shown in the top panel. The dashed red line corresponds to the Doppler shift induced by the planet’s orbital motion. The alignment of this trail with the predicted velocity curve, derived from WASP-33 b’s orbital parameters, confirms the detection of the planet’s atmospheric signal. The cross-correlation map (bottom panel in Fig. 5) shows the composite atmospheric thermal emission with SNR = 7.4. The dashed white lines indicate the expected planet’s radial velocity () from F23 and a system velocity () from Gontcharov (2006). SNR is calculated following the same approach outlined in Kasper et al. (2021) using the 3 clipping method. The positive cross-correlation value further suggests that the detected spectral lines are in emission rather than absorption, supporting the presence of a thermally inverted dayside atmosphere.
@@
We also attempted to detect individual molecular signals from CO, H2O, and OH in the atmosphere using models in Figure 4. These results are shown in Figure 6, formatted similarly to the result of the composite atmospheric model. CO is robustly detected at the expected planetary velocity, consistent with previous studies and confirming its presence in the planet’s dayside emission. With the most significant detection of CO, we tested the impact of the 71st order and the order selection where the CO opacity is dominant to improve the detection significance following the approach of Kanumalla et al. (2024). Using the full order set (53 orders) applied for the other detections resulted in a CCF SNR of 5.20, while removing the 71st order yielded 5.03. By selecting only the orders where the CO opacity is dominant (21 orders; see Figure 4), we yielded a CCF SNR of 6.30; without the 71st order, this decreased to 6.17. Therefore, we selected the 21-order set and validated that the impact of the 71st order on the detection significance is approximately 2–3%. The cross-correlation results of H2O and OH exhibit velocity shifts, which may suggest additional atmospheric processes at play, as individual molecular species have shown similar shifts in previous studies (Brogi et al., 2023; Cont et al., 2021, 2024). One possibility is that the intense dayside heating leads to the thermal dissociation of H2O into H and OH, resulting in non-uniform molecular distributions across the atmosphere (e.g., Lothringer et al., 2018; Parmentier et al., 2018; Nugroho et al., 2021, Finnerty et al., in review). Alternatively, strong atmospheric dynamics, such as high-altitude winds or circulation patterns, may contribute to the observed Doppler shifts by redistributing molecular species across different regions of the atmosphere. Follow-up observations, particularly those covering a broader orbital phase range, will be essential to confirm these velocity shifts and further investigate the planet’s atmospheric circulation and chemistry.
Notably, we achieved high-SNR molecular detections of the planetary signal (covering only 0.05 orbital phases) with just 1.46 hours of observing time. While a direct, quantitative comparison with previous studies (e.g., Nugroho et al., 2021; Finnerty et al., 2023) is not appropriate due to differences in instruments and observational setups, those studies typically required longer observation times of 2–4 hours to reach comparable detection significance. Thus, our result qualitatively demonstrates the sensitivity and efficiency of IGRINS-2 for HRCCS.